What is the percent of carbon in a pure sample of glucose?
1
Expert's answer
2020-04-20T15:09:11-0400

6 moles of carbon will have 12.01g/mol x 6 = 72.06g of Carbon, 12 moles of Hydrogen will have the mass of 12 x 1.008 = 12.096g of Hydrogen, 6 moles of oxygen will have 16.00 x 6 = 96 g of Oxygen. Thus, 1 mole of Glucose (C6H12O6) has a total mass of 72.06 +12.096 + 96 = 180.16 g/mol

Mass fraction of Carbon = 72.06g/180.16g = 0.4000

Therefore, mass % of Carbon= 0.4000×100 = 40.00%